OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (jdoe0001):

\(\bf cos(3x)\implies cos(2x+x)\implies cos(2x)cos(x)-sin(2x)sin(x)\) expand the 2x angles, check your trig identites for the "double-angle identities"

OpenStudy (anonymous):

thanks for actually answering but its the whole double angle thing that I don't understand. like I can expand it and all but from there im lost

OpenStudy (jdoe0001):

ok, so \(\bf cos(3x)\implies cos(2x+x)\implies cos(2x)cos(x)-sin(2x)sin(x)\\ \quad \\ \textit{let us use the cos(2x) of }\quad 1-2sin^2(x)\qquad thus\\ \quad \\\ [\quad 1-2sin^2(x)\quad ]cos(x)-[\quad 2sin(x)cos(x)\quad ]sin(x)\\ \quad \\\ [cos(x)-2sin^2(x)cos(x)]-[2sin^2(x)cos(x)]\implies cos(x)-4sin^2(x)cos(x)\\ \quad \\ cos(x)[1-4sin^2(x)]\implies cos(x)[1^2-4sin^2(x)]\\ \quad \\ \textit{recall that}\qquad a^2-b^2 = (a-b)(a+b)\\ \quad \\ cos(x)[1^2-2^2sin^2(x)]\implies cos(x)[1^2-(2sin(x))^2]\\ \quad \\ cos(x)[1^2-(2sin(x))^2]\implies cos(x)[1-2sin(x)][1+2sin(x)]\)

OpenStudy (jdoe0001):

hmm got a bit truncated \(\bf cos(3x)\implies cos(2x+x)\implies cos(2x)cos(x)-sin(2x)sin(x)\\ \quad \\ \textit{let us use the cos(2x) of }\quad 1-2sin^2(x)\qquad thus\\ \quad \\\ [\quad 1-2sin^2(x)\quad ]cos(x)-[\quad 2sin(x)cos(x)\quad ]sin(x)\\ \quad \\\ [cos(x)-2sin^2(x)cos(x)]-[2sin^2(x)cos(x)]\\ \quad \\\implies cos(x)-4sin^2(x)cos(x)\\ \quad \\ cos(x)[1-4sin^2(x)]\implies cos(x)[1^2-4sin^2(x)]\\ \quad \\ \textit{recall that}\qquad a^2-b^2 = (a-b)(a+b)\\ \quad \\ cos(x)[1^2-2^2sin^2(x)]\implies cos(x)[1^2-(2sin(x))^2]\\ \quad \\ cos(x)[1^2-(2sin(x))^2]\implies cos(x)[1-2sin(x)][1+2sin(x)]\)

OpenStudy (anonymous):

is there any way you can break that down into English? O.o

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

\(\bf cos(3x)\implies cos(2x+x)\implies cos(2x)cos(x)-sin(2x)sin(x)\\ \quad \\ \color{red}{\textit{let us use the cos(2x) of }\quad 1-2sin^2(x)\qquad thus}\\ \quad \\\ [\quad 1-2sin^2(x)\quad ]cos(x)-[\quad 2sin(x)cos(x)\quad ]sin(x)\\ \quad \\\ [cos(x)-2sin^2(x)cos(x)]-[2sin^2(x)cos(x)]\\ \quad \\\implies cos(x)-4sin^2(x)cos(x)\\ \quad \\ cos(x)[1-4sin^2(x)]\implies cos(x)[1^2-4sin^2(x)]\\ \quad \\ \color{red}{\textit{recall that}\qquad a^2-b^2 = (a-b)(a+b)}\\ \quad \\ cos(x)[1^2-2^2sin^2(x)]\implies cos(x)[1^2-(2sin(x))^2]\\ \quad \\ cos(x)[1^2-(2sin(x))^2]\implies cos(x)[1-2sin(x)][1+2sin(x)]\)

OpenStudy (jdoe0001):

is pretty straightforward, we're just using the \(1-2sin^2(x)\) identity for the cos(2x) then we get 2 values added that give us a difference of squares, which we then expand into 2 binomials

OpenStudy (anonymous):

ohhhh.....I think I get it a little better now. thanks!

OpenStudy (jdoe0001):

yw